Find the common factors of the given terms. 
12x, 36 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Factors of 12x and 36
12x = 2×2×3×x
36 = 2×2×3×3
Common factors of 12x and 36 are 2, 2, 3
and , 2×2×3 = 12

Thanks